A ship is 3.6km from a lighthouse on a bearing of 215 degrees. How far west (to the nearest tenth of a kilometre) is the ship from the lighthouse?
I am so bad at these bearing questions! I really don't get what to do. Is it possible if you could link some websites about bearings? I tried to find some but I don't know which one is relevant to my question. Thank you in advance for any help!!

Answer by rothauserc(4718)

bearings are measured clockwise from the north line
:
a bearing of 215 degrees is 215 - 180 = 35 degrees with a hypotenuse of length 3.6km
:
we want the horizontal distance which is one side of the right triangle formed by the bearing, south line and the horizontal line to the bearing
:
let h be the horizontal distance
:
sine(35) = h / 3.6
:
h = sine(35) * 3.6 = 2.064875171 approx 2.1
:
**********************************
the ship is 2.1 km west
**********************************
:
google ship on a bearing
: